OpenStudy (anonymous):

3 questions will fan an medal the best answer What is the constant of variation for the quadratic variation? x 2 3 4 5 6 y 32 72 128 200 288 A. 8 B. 16 C. 24 D. 40 What is the constant of variation for the quadratic variation? 9y = 4x2 A.-2 B.4/9 C.2/3 D.9/4 If f(x) varies directly with x2, and f(x) = 75 when x = 5, find the value of f(4). A. 100 B. 50 C. 48 D. 25 Write a quadratic variation equation

OpenStudy (anonymous):

Write a quadratic variation equation if g(x) varies directly with x2, and g(x) = 98 when x = 7. A. g(x) =1/2 x2 B. g(x) = 2x2 C. g(x) = 14x2 D. g(x) = 7x2

OpenStudy (anonymous):

agl202?

OpenStudy (anonymous):

For #1, use the model y=a*x^2. Let x be 2, and y be 32.

OpenStudy (anonymous):

Please bring out pencil and paper and find a.